Consider the combustion of propane:

C3H8 (g) + 5O2 (g) → 3CO2 (g) + 4H2O(l) ΔH = –2221 kJ

Assume that all of the heat comes from the combustion of propane. Calculate ΔH in which 5.00 g of propane is burned in excess oxygen at constant pressure.
Answer #1

The reaction of combustion of propane is

C3H8 (g) + 5O2 (g) → 3CO2 (g) + 4H2O(l)        ΔH = –2221 kJ

The enthalpy of combustion of 1 mole of propane is -2221 kJ.

Number of moles of propane in 5 g of propane = 5/44.1 = 0.113 mol

Therefore, dH for 5 g (0.113 mol) of propane = –2221 x 0.113 kJ = - 251.81 kJ